In a certain Poisson frequency distribution the frequency corresponding to 3 successes is
one-third the frequency corresponding to 4 successes. Find its mean and standard deviation.
1
Expert's answer
2019-04-24T11:38:04-0400

"P(X=3)=\frac{1}{3}P(X=4)\\
e^{-\lambda}\frac{\lambda^3}{3!}=\frac{1}{3}e^{-\lambda}\frac{\lambda^4}{4!}\\
\lambda=\frac{3*4!}{3!}=12\\
Mean=Variance=\lambda=12\\
Standard\; deviation=\sqrt{12}=2\sqrt{3}"